OpenStudy (anonymous):

f(x)=1 , f(x+h)-f(h) h doesnot equal 0 ___x _________ 2 h

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh i see \[f(x)=\frac{x}{2}\] maybe ? and you want \[\frac{f(x+h)-f(x)}{h}\] is that it?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

except that the second part is f(h) not f(x)

OpenStudy (anonymous):

that is a typo i bet

OpenStudy (anonymous):

honestly, i wouldnt know....thats whats on my packet lol

OpenStudy (anonymous):

wait your right just looked at it sorry

OpenStudy (anonymous):

of course i am right. this is the "difference quotient." you write \[\frac{\frac{x+h}{2}-\frac{x}{2}}{h}\]

OpenStudy (anonymous):

then some algebra to get \[\frac{x+h-x}{2h}\] and finally just \[\frac{1}{2}\] i.e the slope